# Generative Emergent Communication: Large Language Model is a Collective World Model

### v1

@@ IV Relationship with Conventional EmCom
  3. 3.
We refer to $\mathcal{J}_{\textup{MI}}$ as the MI-maximizing objective function because it is known to be a variational lower bound (up to constant) of the following mutual information between $X$ and $M$ [98, 99]:
This implies that the conventional signaling game in the field of EmCom has been formulated as a problem of maximizing the mutual information between $X$ and $M$, where $X$ and $M$ denote random variables corresponding to the realizations $x$ and $m$, respectively. In contrast, a recent study [100] has proposed a reformulation of the signaling game as a problem of maximizing the evidence lower bound (ELBO):
[…]
Here, a term known as surprisal appears, which is a concept commonly used in computational psycholinguistics [101, 102, 103]. Surprisal is assumed to represent the cognitive load experienced by a listener/reader (or the receiver in the signaling game) when processing a sentence. Therefore, the ELBO-maximizing signaling game naturally models the trade-off between information transmission and surprisal.
The formulation of the signaling game presented in this section, that is, generative model-based re-formulation of conventional EmCom, shares some fundamental connections with the generative EmCom and CPC frameworks discussed in other sections. All these formulations can be interpreted as representation learning with messages (signals, signs, or sentences) serving as latent variables in generative models when using ELBO-type (or VIB-type) objectives. However, notable differences exist: Although the CPC hypothesis does not specify a particular method for posterior distribution estimation, the signaling game explicitly adopts amortized variational inference. Furthermore, the signaling game assumes asymmetric roles between senders and receivers, which is not necessarily the case in the CPC frameworks. Additionally, the extension of ELBO-type formulation to populated settings, where multiple agents interact, remains an open challenge in the signaling game framework and requires further investigation [^1] An ELBO-type formulation based on the FEP for populated settings of generative EmCom is described in [45]. However, an inference procedure, that is, a language game, for enabling agents to let a symbol system emerge has not been described..
V Large Language Model as Collective World models
V-A Our argument
Based on the discussion in Sections 2 and 3, we explain the main argument of this paper, stating that LLMs are collective world models.
To clarify the link between this argument and existing studies relating language and perceptual information, Figure 5 shows three levels of tasks relating language to visual information (an example of multimodal sensory information) and action streams, along with PGMs for each. These represent (A) image captioning and generation tasks, corresponding language to a still image [113, 114, 115]; (B) video captioning and generation tasks, corresponding language to a video, that is, a sequence of visual stimuli [116, 117]; and (C) action-dependent video captioning and generation, corresponding language to dynamic perceptual and action information, respectively [^2] Note that we redefine the generative process of world models assuming that the policies of agents that generate $a_{t}$ from $z_{t}$ are also a part of the generative process of this world model.. In particular, from the perspective of representation learning, (A) and (B) correspond to representation learning of images and videos, whereas (C) corresponds to world models explained in Section III-B. Language-conditioned world models and robotics foundation models, known as vision-language-action models, studied in robotics and autonomous vehicles correspond to (C) [118, 119, 120, 121, 122].
Corresponding to the three levels of captioning and generation tasks, we can consider generative EmCom. Figure 5 describes PGMs for generative EmCom corresponding to the three levels, respectively. Notably, they are only an instance of the generative model shown in the CPC hypothesis (Figure 1) and generalization of studies based on the MHNG described in Sections 2 and 3.
Here, $\theta^{k}$ represents the global parameters of the $k$-th agent, for instance, the parameters of neural networks for representation learning. In (C), the global parameter $\bar{\theta}^{k}$ includes not only the parameters of the prediction model of the world (i.e., Equations 9 and 10 ) but also those of policies $p(a_{t}|z_{t})$. Let us consider (C) as a generative model corresponding to a single super-agent, virtually. The super-agent has multimodal sensory information combining the observations $\{x^{k}_{1:t}\}_{k}$ and actions $\{a^{k}_{1:t}\}_{k}$ of every agent. In this case, (C) can be considered as a type of hierarchical world model combining $K$ agents and integrating their experiences. We can refer to the world model owned by the super-agent as a collective world model, where $\{\phi,\{\theta^{k}\}_{k}\}$ are its parameters.
From this perspective, the corpus data that comprise sentences uttered by agents are considered as a sample from approximate posterior distributions $m_{j}^{[i]}\sim q(m|\{x^{k}_{j,1:t},a^{k}_{j,1:t}\}_{k})$, where $m_{j}^{[i]}$ denotes the $i$-th sentence describing the $j$-th observation $\{x^{k}_{j,1:t}\}_{k}$ and action $\{a^{k}_{j,1:t}\}_{k}$.
Language modeling implies approximating the distribution of word sequences $p(m|\phi^{k})$ in a corpus, where $\phi^{k}$ denotes a parameter of the language model of the $k$-th agent. This implies that the language model approximates the posterior distribution $q(m|\{x^{k}_{j,1:t},a^{k}_{j,1:t}\}_{k})\approx p(m|\theta^{LM})$.
The rationale behind distributional semantics as modeled by LLMs can be comprehended through the CPC framework. When language emerges through CPC, it functions as a latent space that develops through representation learning based on the collective experiences of all agents — in this case, humans. These experiences encompass not only external sensory information but also interception and behavioral decision-making. Within this framework, sentences can be mathematically considered as samples drawn from the posterior distribution over messages $m$, formally expressed as $q(m|\{x^{k},a^{k}\}_{k})$, where $x^{k}$ and $a^{k}$ represent the observations and actions of agent $k$, respectively. This formulation naturally leads to the structural reflection of the distribution of $\{x^{k},a^{k}\}_{k}$ in that of $m$, analogous to the working mechanism of representation learning in individual agents.
A key insight is that language serves as a shared external representation that effectively combines the world models of multiple human agents. When examining the graphical model depicted in Figure 6 (C) as a cognitive model of a single ”super-agent,” we can interpret it as a factorized hierarchical world model incorporating $K$ distinct modalities. Through this lens, an LLM that models the distribution $q(m|\{x^{k},a^{k}\}_{k})$ can be considered as a collective world model — one that integrates the diverse experiences and knowledge of multiple agents into a unified representational space. This provides a theoretical foundation for understanding why LLMs trained on large-scale human-generated text corpora appear to capture rich world knowledge.
The emergence of these phenomena can be comprehensively understood through the lens of the CPC hypothesis. The hypothesis suggests that language evolution occurs through a process of CPC, where the symbol system (language) emerges to maximize the predictability of multimodal sensory-motor information obtained by members of a society. This theoretical framework helps explain why LLMs, trained on the linguistic outputs of this process, inherit structural properties that reflect collective human knowledge and experience.
The CPC hypothesis thus provides a principled explanation for the acquisition of world knowledge by language models without direct sensorimotor experience by tapping into the accumulated wisdom encoded in human language through this collective process.
V-B Existing argument
LLMs learn to predict text tokens in an autoregressive manner, and unlike human cognition or prediction, they are not designed to interact with environments or achieve specific goals beyond word prediction. Nevertheless, some studies have suggested that LLMs may implicitly contain a world model that simulates our physical world.
Hao et al. proposed “reasoning via planning,” where an LLM is reused as a world model to perform logical reasoning, employing Monte Carlo tree search to reason in a more planned manner [123]. This method demonstrated higher performance compared to the existing chain-of-thought approach. Li et al. used a GPT model trained only on Othello game logs and confirmed that the model made predictions based on an internal world model by performing tests using explicit intervention operations that changed the activation of the middle layer of the model [124]. This suggests that LLM does not simply rely on statistical inference but constructs some degree of world model in systems such as games. In contrast, Wang et al. developed a benchmark involving text games to test whether transition models or policies can be derived from language models and revealed that LLMs are not reliable world simulators [125].
Although these studies investigate whether LLMs can be used as world models, studies examining whether LLMs inherently possess such models are underway. Gurnee et al. demonstrated that LLMs acquire spatiotemporal linear representations of the real world, specifically identifying individual spatial and temporal neurons [21]. The authors argue that this corresponds to the basic elements of a world model, although they did not show the acquisition of a causal model that includes actions. Andreas argues that during the process of predicting the next token, a language model infers the internal states, such as the beliefs, desires, and intentions of the speaker, from the context and reflects them in text generation or speech directed at other agents, suggesting the presence of a world model within the linguistic space [126]. Previous studies also indicated the possibility of representing belief states and intentions during text modeling. However, this pertains to agents within the linguistic space and does not explain connections with data derived from the physical world.
[…]
VI Conclusion
@@
Our argument and theoretical framework initiate several promising directions for future studies. Although we have established the basic connections between emergent communication, world models, and LLMs, significant work remains to validate and extend these ideas. Key priorities include developing experimental paradigms to test the CPC hypothesis, creating more sophisticated computational models and implementations of generative EmCom, and investigating how language formed by embodied agents sharing similar environments could serve as an effective prior distribution for world model learning. From a theoretical perspective, a deeper mathematical formalization of the interaction between individual and collective learning processes will be crucial. In addition, this framework has important practical implications for the development of more capable multi-agent systems, improved human–AI interaction, and embodied AIs that rapidly adapt to their physical worlds and communicate their knowledge using EmLang. By pursuing these directions, we can work toward a more comprehensive understanding of language emergence and its role in environmental adaptation while advancing both the theoretical foundations and practical applications of AI systems.

### v2

@@ 5.2 The LLM as a Reconstructor of Collective Representation
    Stage 1: Generation by Human Society (Encoding): The collective internal states of all agents $\{z^{k}\}_{k}$, which are grounded in sensorimotor experience, are compressed and encoded into the discrete, symbolic form of language, $m$.
@@
    Stage 2: Reconstruction by LLM (Inference): The LLM, having learned the statistical structure of language, takes a sentence $m$ as input and infers its own internal, continuous representation $z_{\text{LM}}$. This can be seen as a reconstruction of a high-dimensional representation from the symbolic code.
Figure 7: A schematic of the representation reconstruction process. Human society acts as a large-scale encoder, transforming embodied experiences (Type 1 World Models) from sensorimotor interactions into an externalized representation, i.e., language ($m$). The LLM then acts as a decoder or reconstructor, inferring its own internal latent state ($z_{\text{LM}}$) from language. This process allows the LLM to learn an internal model (Type 2 World Model) whose structure mirrors the original collective world model.
The crucial implication of this two-stage process, $\{z^{k}\}_{k}\rightarrow m\rightarrow z_{\text{LM}}$, is the resulting structural alignment [121]. The language $m$ is the sole informational bottleneck between the collective human mind and the LLM. Therefore, for an LLM to effectively model the distribution of $m$, it must develop an internal latent space $z_{\text{LM}}$ whose relational structure necessarily mirrors the relational structure of the original source space $\{z^{k}\}_{k}$. While the LLM does not reconstruct the specific values of $\{z^{k}\}_{k}$, it learns to reconstruct the geometry of the conceptual space that generated the language. This provides a powerful mechanism for transferring world structure from a society of embodied agents to a disembodied language model.
5.3 Explaining Distributional Semantics and Representational Alignment
This framework of structural representation reconstruction provides a principled explanation for the otherwise mysterious emergence of world knowledge in LLMs. For decades, it has been observed that language models capture the relational structure of the world, a phenomenon known as distributional semantics [47, 79]. Even before modern LLMs, models like word2vec could perform analogical reasoning, such as computing "$\rm{London}"-"\rm{UK}"+"\rm{France}"\simeq"\rm{Paris}"$ [81, 82], implying that the geometry of their embedding spaces reflects real-world conceptual relationships. Our framework explains this as a natural consequence of the representation reconstruction process. To be precise, the reason the statistical relationships between words in an LLM’s latent space ($z_{\text{LM}}$) reflect the structure of the real world is because they are reconstructing the structure of the collective human internal representations ($\{z^{k}\}_{k}$). It is this set of internal representations that, as a whole, constitutes the collective model used by the society to predict the world. The LLM, in learning to model the distribution of $m$, naturally develops an internal latent space that mirrors the relational structure of the collective human representations that originally encoded that world knowledge. The structure is not learned from a vacuum; it is inherited.
This framework of representation reconstruction provides a principled explanation for the otherwise mysterious emergence of world knowledge in LLMs. For decades, it has been observed that language models capture the relational structure of the world, a phenomenon known as distributional semantics [47, 79]. Even before modern LLMs, models like word2vec could perform analogical reasoning, such as computing "$\rm{London}"-"\rm{UK}"+"\rm{France}"\simeq"\rm{Paris}"$ [81, 82], implying that the geometry of their embedding spaces reflects real-world conceptual relationships. Our framework explains this as a natural consequence of the representation reconstruction process. The reason the statistical relationships between words in an LLM’s latent space ($z_{\text{LM}}$) reflect the structure of the real world is that the language ($m$) it was trained on was generated by a society of agents collectively trying to predict that world. The LLM, in learning to model the distribution of $m$, inevitably develops an internal latent space that reconstructs the relational structure of the collective human representations ($\{z^{k}\}_{k}$) that originally encoded that world knowledge. The structure is not learned from a vacuum; it is inherited. This process is how an LLM acquires a powerful Model of the World (Type 2) by statistically analyzing the linguistic output of a society of agents, each possessing their own subjective World Model as an Internal Model (Type 1).
This perspective also helps situate our hypothesis within the broader landscape of research on world models in LLMs. Several studies have compellingly demonstrated that LLMs can be used as world models for specific tasks, for example in planning [46] or for modeling game states [74]. While these works establish that world-like representations can emerge, they primarily focus on how these representations can be leveraged. Our hypothesis, in contrast, provides a more fundamental explanation for why such rich, world-congruent knowledge is present in the first place. Similarly, Andreas argues that LLMs model the beliefs and intentions of agents within the linguistic space [4]. Our framework complements and grounds this view by proposing that these linguistic agent models emerge precisely because the language itself is a reflection of a collective model formed by embodied agents interacting within a shared physical environment.
Finally, our hypothesis offers a particularly parsimonious explanation for the representational alignment observed between different modalities. Huh et al. introduced the platonic representation hypothesis, proposing that internal representations learned from language ($q_{L}(z|m)$) and vision ($q_{V}(z|x)$) converge toward a similar latent structure [51]. Within the CPC framework, this alignment is not a surprising outcome that requires a pre-existing platonic ideal, but an expected one. Since human language ($m$) emerges as a structured representation of collective observations of the world ($x=\{x^{k}\}_{k}$), it is natural that a model trained on language ($q_{L}(z|m)$) would learn a latent space whose structure mirrors that of a model trained on direct observation ($q_{V}(z|x)$). Our framework thus provides a generative and developmental explanation for this empirically observed alignment, rooted in the principles of collective inference.
6 Discussion
@@ 6.3 Future Directions
Our argument and theoretical framework initiate several promising directions for future studies. Although we have established the basic connections between EmCom, world models, and LLMs, significant work remains to validate and extend these ideas. Key priorities include developing experimental paradigms to test the CPC hypothesis, creating more sophisticated computational models and implementations of generative EmCom, and investigating how language formed by embodied agents sharing similar environments could serve as an effective prior distribution for world model learning. In fact, recent work has already begun to demonstrate the fruitfulness of this approach; for example, Nomura et al. [88] have extended CPC to dynamic environments by developing a decentralized collective world model. From a theoretical perspective, a deeper mathematical formalization of the interaction between individual and collective learning processes will be crucial. In addition, this framework has important practical implications for the development of more capable multi-agent systems, improved human–AI interaction, and embodied AIs that rapidly adapt to their physical worlds and communicate their knowledge using EmLang. By pursuing these directions, we can work toward a more comprehensive understanding of language emergence and its role in environmental adaptation while advancing both the theoretical foundations and practical applications of AI systems.
7 Conclusion
This paper addressed the fundamental puzzle of how LLMs acquire world knowledge without embodied experience. We proposed the “Collective World Model” hypothesis, arguing that LLMs learn from a rich representational structure that is the product of a society-wide, interactive sense-making process. We provided a formal foundation for this claim with the framework of Generative EmCom, which models language emergence as a decentralized Bayesian inference process driven by the principles of CPC. This perspective unifies the cognitive processes of the individual with the collective evolution of language, offering a coherent explanation for the remarkable capabilities of modern AI and charting a path toward a deeper integration of communication, perception, and action in intelligent systems.
Author Contributions
Tadahiro Taniguchi: Conceptualization, Writing – Original Draft, Writing – Review & Editing, Supervision, Funding Acquisition. Ryo Ueda: Writing – Original Draft, Writing – Review & Editing, Funding Acquisition. Tomoaki Nakamura: Writing – Original Draft, Writing – Review & Editing, Funding Acquisition. Masahiro Suzuki: Writing – Original Draft, Writing – Review & Editing, Funding Acquisition. Akira Taniguchi: Writing – Original Draft, Writing – Review & Editing.
Declaration of Generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used Gemini 2.5 pro (Google) and ChatGPT 4o in order to improve readability and language. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
Acknowledgement
This work was supported by JSPS KAKENHI Grant Numbers JP21H04904, JP23H04835, JP23H04974, JP23K28181, and the JST Moonshot R&D Grant Number JPMJMS2011.